# Safety Controller Synthesis for Stochastic Networked Systems under Communication Constraints

### v1

@@ Abstract
This paper develops a framework for synthesizing safety controllers for discrete-time stochastic linear control systems (dt-SLS) operating under communication imperfections. The control unit is remote and communicates with the sensor and actuator through an imperfect wireless network. We consider a constant delay in the sensor-to-controller channel (uplink), and data loss in both sensor-to-controller and controller-to-actuator (downlink) channels. In our proposed scheme, data loss in each channel is modeled as an independent Bernoulli-distributed random process. To systematically handle the uplink delay, we first introduce an augmented discrete-time stochastic linear system (dt-ASLS) by concatenating all states and control inputs that sufficiently represent the state-input evolution of the original dt-SLS under the delay and packet loss constraints. We then leverage control barrier certificates (CBCs) for dt-ASLS to synthesize a controller that guarantees dt-SLS safety in a stochastic sense, ensuring that all trajectories of dt-SLS remain within safe regions with a quantified probabilistic bound. Our approach translates safety constraints into matrix inequalities, leading to an optimization problem that eventually quantifies the probability of satisfying the safety specification in the presence of communication imperfections. We validate our results on an RLC circuit subject to both constant delay and probabilistic data loss.
@@ I Introduction
Motivation. Safety analysis in stochastic networked control systems has become a crucial design aspect due to the increasing application of such systems in safety-critical domains such as autonomous vehicles, networked robotics, and industrial automation. In these contexts, closed-loop controllers rely on wireless communication channels that often introduce effects such as delays and data loss. Even a fixed delay in the feedback loop can degrade performance or even destabilize a system, while data loss in sensor or actuator signals can further disrupt the closed-loop control performance [1]. Designing controllers that formally ensure a networked system remains within safe operational bounds under communication issues—such as delays and stochastic uncertainties—is challenging. The key difficulty lies in maintaining safety despite outdated state information and data loss.
To address the challenges of safe controller synthesis for stochastic control systems, the existing literature has primarily focused on the promising approach of control barrier certificates (CBCs), initially introduced in [2]. Similar to Lyapunov functions, CBCs are designed to satisfy specific conditions on both the function and its evolution along the system trajectories. By defining an initial level set of CBCs from a given set of initial states, these CBCs effectively separate unsafe regions from all admissible system evolutions, thereby providing (probabilistic) safety guarantees. This technique is widely applied for the formal verification and synthesis of both deterministic [3, 4, 5] and stochastic dynamical systems [6, 7, 8, 9, 10, 11, 12].
@@
While CBCs have been effectively used to enforce safety constraints in stochastic control systems, they typically assume a continuously closed feedback loop. However, in the simultaneous presence of delays and data loss, the loop may be intermittently broken—control signals might fail to reach the actuators on time, and control actions may be stalled due to missing sensor measurements. Although several studies have addressed the safety analysis of control systems with time delay [13, 14, 15], they typically assume a reliable feedback loop—ignoring data loss in communication channels. As a result, these methods are inadequate for scenarios where both delay and data loss jointly impact control synthesis and closed-loop behavior.
The challenges of control systems operating over wireless communication channels—where data loss and delays are prevalent—have been investigated in the literature (see e.g.[16, 17, 18]). Approaches like the Glossy architecture[19] improve reliability via synchronous flooding, while studies on Kalman filtering with intermittent observations [20] show that even minor communication failures can degrade control accuracy. Optimal control strategies have been developed to address these issues, including linear quadratic Gaussian (LQG) control under delay-dependent constraints [21] and models incorporating traffic-correlated delays and losses [22]. Broader surveys are available in [23], and recent work has also examined stability under repeated deadline misses [24], emphasizing the need for timeliness in ensuring robustness. The stability analysis of control systems over wireless networks has also been addressed in [25, 26], where the augmentation approach used in our framework is inspired by [25]. However, none of the above studies have considered the coupled effects of delay and data loss in the context of safety analysis for control systems over wireless networks, as we addressed within this work.
Key Contribution. Inspired by the above challenges, we develop a formal framework for synthesizing safe controllers for dt-SLS under communication constraints. Specifically, we consider delays in the sensor-to-controller (uplink) channel and probabilistic data loss in both the uplink and controller-to-actuator (downlink) channels (see Fig. 1). Our approach handles communication imperfections by introducing an augmented system representation (dt-ASLS) that captures the effects of uplink delays and stochastic data loss. We leverage CBCs within this augmented framework to provide probabilistic safety guarantees by formulating safety constraints as matrix inequalities. Unlike our prior studies [27, 28], which assume no communication delay, our proposed framework accounts for the coupled effects of uplink delays and data loss, making the problem significantly more challenging.
@@ II-A Notation and Preliminaries
We denote the sets of real, positive real, and non-negative real numbers, respectively, by $\mathbb{R}$, $\mathbb{R}^{+}$, and $\mathbb{R}_{0}^{+}$. Sets of non-negative and positive integers are denoted by $\mathbb{N}$ and $\mathbb{N}^{+}$. Given $N$ vectors $x_{i}\in\mathbb{R}^{n_{i}}$, $x\!=\![x_{1};\dots;x_{N}]$ is the concatenated column vector. Trace of a matrix $A\in\mathbb{R}^{N\times N}$ with diagonal elements $a_{1},\dots,a_{N}$ is written by $\text{Tr}(A)\!=\!\sum_{i=1}^{N}a_{i}$. An identity matrix of dimension $n$, and a zero matrix of dimension $n\times m$ are denoted, respectively, by $\mathbf{I}_{n}$ and $\mathbf{0}_{n\times m}$. A normal distribution with the mean value $\mu$ and the variance $\sigma$ is denoted by $\mathcal{N}(\mu,\sigma)$. For a symmetric matrix $\mathcal{P}$, $\mathcal{P}\succ 0$ ($\mathcal{P}\succeq 0$) denotes $\mathcal{P}$ is positive definite (positive semi-definite). The transpose of a matrix $A$ is written as $A^{\top}$. The Euclidean norm of a vector $x\in\mathbb{R}^{n}$ is expressed as $\|x\|$. The empty set is denoted by $\emptyset$. A diagonal matrix in ${\mathbb{R}}^{n\times{n}}$ with diagonal entries $(a_{1},\ldots,a_{n})$ is signified by $\mathsf{diag}(a_{1},\ldots,a_{n})$. By $A\Leftrightarrow B$, we denote ‘$A$ if and only if $B$’. The variance of a random variable $a$ is denoted by $\textsf{Var}[a]$. Given a system $\Sigma$ and a property $\Upsilon$, $\Sigma\models\Upsilon$ denotes $\Sigma$ satisfies $\Upsilon$.
By $(\Omega,\mathcal{F}_{\Omega},\mathds{P}_{\Omega})$, we present a probability space of events, with the sample space $\Omega$, $\mathcal{F}_{\Omega}$ being a sigma-algebra on $\Omega$, and a probability measure $\mathds{P}_{\Omega}$. We assume random variables $\mathcal{X}$ are measurable functions, i.e., $\mathcal{X}:(\Omega,\mathcal{F}_{\Omega})\rightarrow(S_{\mathcal{X}},\mathcal{F}_{\mathcal{X}})$, such that $\mathcal{X}$ induces a probability measure on $(S_{\mathcal{X}},\mathcal{F}_{\mathcal{X}})$ as $Prob\{\mathcal{A}\}=\mathds{P}_{\Omega}\{\mathcal{X}^{-1}(\mathcal{A})\}$, $\forall\mathcal{A}\in\mathcal{F}_{\mathcal{X}}$. A topological space $\mathbf{S}$ is said to be Borel, denoted by $\mathcal{B}(\mathbf{S})$, if it is homeomorphic to a Borel subset of a Polish space, i.e., a separable and metrizable space.
@@ II-B Networked Control System
The collocated control and estimation unit of the system is assumed to be remote, communicating with the sensing and actuator stations via wireless communication channels. We assume the communication channel introduces both delay and data loss during data transmission. To account for communication issues, we introduce a constant delay $\tau\in\mathbb{N}$ in the uplink channel, i.e., between the sensor and the controller. Additionally, to model data loss in the uplink and downlink channels, we consider two mutually independent random processes $\theta_{k}$ and $\phi_{k}$, $k\in\mathbb{N}$, where the corresponding binary random variables are selected from Bernoulli distributions. Specifically, $\theta_{k}=1$ indicates that the state information sent to the controller at time $k$ is successfully received, while $\theta_{k}=0$ denotes a data loss at that time. Note that due to the uplink delay, if a data packet is not dropped, it is received by the controller after $\tau$ time steps. Similarly, $\phi_{k}=1$ indicates that the actuator successfully receives the control input at time $k$, while $\phi_{k}=0$ means the control input is not received at that time. Furthermore, $p_{\theta}$ and $q_{\phi}$ denote, respectively, the probabilities of successful data transmissions in uplink and downlink channels, i.e., $\mathbb{P}[\theta_{k}=1]=p_{\theta}$, and $\mathbb{P}[\phi_{k}=1]=q_{\phi}$, for any $k\in\mathbb{N}$.
@@
The networked system, comprising the dt-SLS plant in (1) and a remote control unit communicating over wireless channels with delay and data loss, can be formally defined by the following tuple:
@@
We represent $\Sigma$ using two state-space models: one for the plant, denoted by $\Sigma_{P}$, and one for the control unit, denoted by $\Sigma_{C}$. The dt-SLS-$\Sigma$ then evolves according to the following dynamics:
@@
where $\mathcal{F}\in\mathbb{R}^{n\times n}$ is a control matrix, $\hat{x}_{k}^{c}\in X$ and $\hat{u}_{k}^{c}\in U$ are the state and input of the controller at time $k$, and $u_{k}$ is the control signal applied to the actuator at time $k$. In fact, $\hat{u}_{k}^{c}$ is the control signal computed at the control unit at time $k$, and according to (3), if $\phi_{k}=1$, then $u_{k}=\hat{u}_{k}^{c}$. The initial state, denoted by $x_{0}\sim\mathcal{G}$ with $x_{0}\in X$, is drawn from a known distribution $\mathcal{G}$ with finite variance. We denote by $\mathsf{x}_{k}$ the solution process of dt-SLS-$\Sigma$ at time $k\in\mathbb{N}$, under input and process noise trajectories $u(\cdot)$ and $w(\cdot)$, starting from an initial state $x_{0}\in X$.
@@
The dynamics in (3) and (4) capture the interaction between the plant and remote controller over unreliable wireless channels. We represent $\Sigma$ into $\Sigma_{P}$ and $\Sigma_{C}$ to separate the plant dynamics from the controller logic, which is affected by delays and data loss. Due to the uplink delay, the control unit always receives delayed state information. Specifically, at time $k$, the most recent state available to the controller is $x_{k-\tau}$, provided the uplink transmission at time $k-\tau$ was successful (i.e., $\theta_{k-\tau}=1$). The control input $u_{k}$ applied to the plant is $\hat{u}_{k}^{c}$ if $\phi_{k}=1$, or the previous input $u_{k-1}$ if $\phi_{k}=0$, i.e., a zero-order hold mechanism is used in the event of downlink data loss (cf. equation (3)). The controller evolves based on an internal state $\hat{x}_{k}^{c}$, representing its estimate of the plant state at time $k$ based on the available information, and applies a state-feedback control input $\mathcal{F}\hat{x}_{k}^{c}$ (cf. (4), (9)).
@@
Note that the controller does not have access to the variables $\phi_{k}$ in (6) since these occur on the actuator side and no feedback on $\phi_{k}$ is provided to the controller. Consequently, when constructing the networked control system model dt-SLS-$\Sigma$ in Definition 1, the controller always assumes that the control input applied to the plant $\Sigma_{P}$ is $\hat{u}_{k}^{c}$ (i.e., $\phi_{k}=1$). Then, we define the controller’s state at time $k$ as $\hat{x}^{c}_{k}=\mathds{E}\big{[}{x}_{k}\,\big{|}\,\phi_{k}=1,\forall\,k\,,\mathcal{I}_{k}\big{]}$, where $\mathcal{I}_{k}$ is the set of available information at the control side at time $k$, defined as
[…]
@@
$+\sum_{t=0}^{\tau-1}A^{t}w_{k-t-1}\Big{)}\!+\underbrace{(1-\theta_{k-\tau})(A\hat{x}^{c}_{k-1}\!+\!B\hat{u}_{k-1}^{c})}_{(\star)}\!\mid\!\mathcal{I}_{k}\Big{]}\!.$
We have formulated ()($\star$)( ⋆ ) based on the information set $\mathcal{I}_{k}$ available to the control unit at time $k$. If $\theta_{k-\tau}=0$, the control unit does not receive the plant state $x_{k-\tau}$, which—due to the $\tau$-step delay—is the most recent state the controller could access at time $k$. In this case, the control unit uses the model in (4) to compute a model-based estimate of the state $\hat{x}_{k}^{c}$, which serves as the state of the controller. Since $w_{k}$ is a sequence of i.i.d. random variables drawn from $\mathcal{N}(0,\sigma_{w})$, and $\mathds{E}[\theta_{k-\tau}]=p_{\theta}$, we can rewrite (7) as
@@
Given a dt-SLS-$\Sigma=(X,U,A,B,w,\tau,\theta,\phi)$  as in Definition 1 subjected to delay and data loss, consider a safety specification $\Upsilon=(X_{0},X_{1},\mathcal{T})$, where $X_{0},X_{1}\subseteq X$ are its initial and unsafe sets, respectively. We assume $X_{0}\cap X_{1}=\emptyset$, otherwise system is unsafe with a probability of 1. The dt-SLS-$\Sigma$ is said to be safe within time horizon $\mathcal{T}\in\mathbb{N}$, denoted by $\Sigma\models\Upsilon$, if all trajectories of $\Sigma$ starting from $X_{0}$ do not reach $X_{1}$ during $\mathcal{T}$. Since trajectories of dt-SLS-$\Sigma$ are probabilistic, we are interested in computing $\mathds{P}\{\Sigma\models\Upsilon\}\geq 1-\xi_{\tau\theta\phi}$, (cf. Remark 2) where $\xi_{\tau\theta\phi}\in[0,1]$.
@@
With all definitions and concepts in place, we are now ready to state the main problem addressed in this work.
Problem 1
Consider a networked control system described by the dt-SLS-$\Sigma=(X,U,A,B,w,\tau,\theta,\phi)$ in Definition 1 in the presence of delay and data loss, with a given safety specification $\Upsilon=(X_{0},X_{1},\mathcal{T})$. Synthesize a formal controller based on CBCs to ensure the satisfaction of safety specification $\Upsilon$ within a time horizon $\mathcal{T}$ with a guaranteed probabilistic bound $\xi_{\tau\theta\phi}\in[0,1]$, i.e.,
$\mathds{P}\Big{\{}\Sigma\models\Upsilon\Big{\}}\geq 1-\xi_{\tau\theta\phi}.$

### v2

@@ Abstract
This paper develops a framework for synthesizing safety controllers for discrete-time stochastic linear control systems (dt-SLS) operating under communication imperfections. The control unit is remote and communicates with the sensor and actuator through an imperfect wireless network. We consider a constant delay in the sensor-to-controller channel (uplink), and data loss in both sensor-to-controller and controller-to-actuator (downlink) channels. In our proposed scheme, data loss in each channel is modeled as an independent Bernoulli-distributed random process. To systematically handle the uplink delay, we first introduce an augmented discrete-time stochastic linear system (dt-ASLS) by concatenating all states and control inputs that sufficiently represent the state-input evolution of the original dt-SLS under the delay and packet loss constraints. We then leverage control barrier certificates for dt-ASLS to synthesize a controller that ensures the stochastic safety of dt-SLS, guaranteeing that all trajectories remain outside unsafe regions with a quantified probabilistic bound. Our approach translates safety constraints into matrix inequalities, leading to an optimization problem that eventually quantifies the probability of satisfying the safety specification in the presence of communication imperfections. We validate our results on an RLC circuit subject to both constant delay and probabilistic data loss.
@@ I Introduction
Motivation. Safety analysis in stochastic networked control systems has become a crucial design aspect due to the increasing application of such systems in safety-critical domains such as autonomous vehicles, networked robotics, and industrial automation. In many such systems, the controller is located away from the plant (cf. Fig. 1) for both practical and safety reasons. Representative examples include process industries, where controllers are housed in dedicated control rooms outside hazardous zones, and offshore or remote assets (e.g., wind turbines, pipelines) controlled from onshore facilities. In these settings, closed-loop operation depends on wireless communication channels, which are inherently subject to imperfections such as delays and data loss. Even a fixed delay in the feedback loop can degrade performance or destabilize a system, while data loss in sensor or actuator signals can further disrupt the closed-loop control performance [1]. Hence, designing controllers that formally ensure a networked system remains within safe operational bounds under communication issues, such as delays and stochastic uncertainties, is challenging.
To address the challenges of safe controller synthesis for stochastic control systems, the existing literature has primarily focused on the promising approach of control barrier certificates (CBCs), initially introduced in [2]. Similar to Lyapunov functions, CBCs are designed to satisfy specific conditions on both the function and its evolution along the system trajectories. By defining an initial level set of CBCs from a given set of initial states, these CBCs effectively separate unsafe regions from all admissible system evolutions, thereby providing (probabilistic) safety guarantees. This technique is widely applied for the formal verification and synthesis of both deterministic [3, 4] and stochastic dynamical systems [5, 6, 7].
@@
While CBCs have been effectively used to enforce safety constraints in stochastic control systems, they typically assume a continuously closed feedback loop. However, in the simultaneous presence of delays and data loss, the loop may be intermittently broken: control signals might fail to reach the actuators on time, and control actions may be stalled due to missing sensor measurements. Although several studies have addressed the safety analysis of control systems with time delay [8, 9, 10], they typically assume a reliable feedback loop, while ignoring data loss in communication channels. As a result, these methods are inadequate for scenarios where both delay and data loss jointly impact control synthesis and closed-loop behavior.
The challenges of control systems operating over wireless communication channels, where data loss and delays are prevalent, have been investigated in the literature (see e.g.[11, 12, 13]). Approaches like the Glossy architecture[14] improve reliability via synchronous flooding, while studies on Kalman filtering with intermittent observations [15] show that even minor communication failures can degrade control accuracy. Optimal control strategies have been developed to address these issues, including linear quadratic Gaussian control under delay-dependent constraints [16] and models incorporating traffic-correlated delays and losses [17]. Broader surveys are available in [18], and the work [19] has also examined stability under repeated deadline misses, emphasizing the need for timeliness in ensuring robustness. The stability analysis of control systems over wireless networks has also been addressed in [20, 21], where the augmentation approach used in our framework is inspired by [20]. Nevertheless, none of the above studies have considered the coupled effects of delay and data loss in the context of safety analysis for control systems over wireless networks, as we addressed within this work.
Key Contribution. Inspired by the above challenges, we develop a formal framework for synthesizing safe controllers for dt-SLS under communication constraints. Specifically, we consider delays in the sensor-to-controller (uplink) channel and probabilistic data loss in both the uplink and controller-to-actuator (downlink) channels (see Fig. 1). Our approach handles communication imperfections by introducing an augmented system representation (dt-ASLS) that captures the effects of uplink delays and stochastic data loss. We leverage CBCs within this augmented framework to provide probabilistic safety guarantees by formulating safety constraints as matrix inequalities. Unlike our prior work [22], which assumes no communication delay, the proposed framework accounts for the coupled effects of uplink delays and data loss, making the problem significantly more challenging.
@@ II-A Notation and Preliminaries
We denote the sets of real, non-negative and positive real numbers, respectively, by $\mathbb{R}$, $\mathbb{R}_{0}^{+}$, and $\mathbb{R}^{+}$. Sets of non-negative and positive integers are denoted by $\mathbb{N}$ and $\mathbb{N}^{+}$, respectively. Given $N$ vectors $_{i}\in\mathbb{R}^{n_{i}}, x=[x$$x_{1};\dots;x_{N}$] is its concatenated column vector. Trace of a matrix $A\in\mathbb{R}^{N\times N}$ with diagonal elements $a_{1},\dots,a_{N}$ is written as $\text{Tr}(A)\!=\!\sum_{i=1}^{N}a_{i}$. An identity matrix of dimension $n$, and a zero matrix of dimension $n\times m$ are denoted, respectively, by $\mathbf{I}_{n}$ and $\mathbf{0}_{n\times m}$. A normal distribution with the mean value $\mu$ and the variance $\sigma$ is denoted by $\mathcal{N}(\mu,\sigma)$. For a symmetric matrix $\mathcal{P}$, $\mathcal{P}\succ 0$ ($\mathcal{P}\succeq 0$) denotes $\mathcal{P}$ is positive definite (positive semi-definite). The transpose of a matrix $A$ is written as $A^{\top}$. The Euclidean norm of a vector $x\in\mathbb{R}^{n}$ is expressed as $\|x\|$. The empty set is denoted by $\emptyset$. A diagonal matrix in ${\mathbb{R}}^{n\times{n}}$ with diagonal entries $(a_{1},\ldots,a_{n})$ is signified by $\mathsf{diag}(a_{1},\ldots,a_{n})$. The variance of a random variable $a$ is denoted by $\textsf{Var}[a]$. Given a system $\Sigma$ and a property $\Upsilon$, $\Sigma\models\Upsilon$ denotes $\Sigma$ satisfies $\Upsilon$.
By $(\Omega,\mathcal{F}_{\Omega},\mathds{P}_{\Omega})$, we denote a probability space, where $\Omega$ is the sample space, $\mathcal{F}_{\Omega}$ is a sigma-algebra on $\Omega$, and $\mathds{P}_{\Omega}$ is the probability measure. We assume random variables $\mathcal{X}$ are measurable functions, i.e., $\mathcal{X}:(\Omega,\mathcal{F}_{\Omega})\rightarrow(S_{\mathcal{X}},\mathcal{F}_{\mathcal{X}})$, such that $\mathcal{X}$ induces a probability measure on $(S_{\mathcal{X}},\mathcal{F}_{\mathcal{X}})$ as $Prob\{\mathcal{A}\}=\mathds{P}_{\Omega}\{\mathcal{X}^{-1}(\mathcal{A})\}$, $\forall\mathcal{A}\in\mathcal{F}_{\mathcal{X}}$. A topological space $\mathbf{S}$ is said to be Borel, denoted by $\mathcal{B}(\mathbf{S})$, if it is homeomorphic to a Borel subset of a Polish space, i.e., a separable and metrizable space.
@@ II-B Networked Control System
The collocated control and estimation unit of the system is assumed to be remote, communicating with the sensing and actuator stations via wireless communication channels. We assume the communication channel introduces both delay and data loss during data transmission. To account for communication issues, we introduce a constant delay $\tau\in\mathbb{N}$ in the uplink channel, i.e., between the sensor and the controller. Additionally, to model data loss in the uplink and downlink channels, we consider two mutually independent random processes $\theta_{k}$ and $\phi_{k}$, $k\in\mathbb{N}$, where the corresponding binary random variables are selected from Bernoulli distributions. Specifically, $\theta_{k}=1$ indicates that the state information sent to the controller at time $k$ is successfully received, while $\theta_{k}=0$ denotes a data loss at that time. Note that due to the uplink delay, if a data packet is not dropped, it is received by the controller after $\tau$ time steps. Similarly, $\phi_{k}=1$ indicates that the actuator successfully receives the control input at time $k$, while $\phi_{k}=0$ means the control input is not received at that time. Furthermore, $p_{\theta}$ and $q_{\phi}$ denote, respectively, the probabilities of successful data transmissions in uplink and downlink channels, i.e., $\mathds{P}[\theta_{k}=1]=p_{\theta}$, and $\mathds{P}[\phi_{k}=1]=q_{\phi}$, for any $k\in\mathbb{N}$.
@@
The networked system, comprising the dt-SLS plant in (1) and a remote control unit communicating over wireless channels with delay and data loss, can be formally defined by the following tuple
@@
We represent $\Sigma$ using two state-space models: one for the plant, denoted by $\Sigma_{P}$, and one for the control unit, represented by $\Sigma_{C}$. The dt-SLS-$\Sigma$ then evolves according to the following dynamics:
@@
where $\mathcal{F}\in\mathbb{R}^{n\times n}$ is a control matrix, $\hat{x}_{k}^{c}\in X$ (the estimate of plant state in (8)) and $\hat{u}_{k}^{c}\in U$ are the state and input of the controller at time $k$, and $u_{k}$ is the control signal applied to the actuator at time $k$. In fact, $\hat{u}_{k}^{c}$ is the control signal computed at the control unit at time $k$, and according to (3), if $\phi_{k}=1$, then $u_{k}=\hat{u}_{k}^{c}$. The initial state, denoted by $x_{0}\sim\mathcal{G}$ with $x_{0}\in X$, is drawn from a known distribution $\mathcal{G}$ with finite variance. We denote by $\mathsf{x}_{k}$ the solution process of dt-SLS-$\Sigma$ at time $k\in\mathbb{N}$, under input and process noise trajectories $u(\cdot)$ and $w(\cdot)$, starting from an initial state $x_{0}\in X$.
@@
The dynamics in (3) and (4) capture the interaction between the plant and remote controller over unreliable wireless channels. We represent $\Sigma$ into $\Sigma_{P}$ and $\Sigma_{C}$ to separate the plant dynamics from the controller logic, which is affected by delays and data loss. Due to the uplink delay, the control unit always receives delayed state information. Specifically, at time $k$, the most recent state available to the controller is $x_{k-\tau}$, provided the uplink transmission at time $k-\tau$ was successful (i.e., $\theta_{k-\tau}=1$). The control input $u_{k}$ applied to the plant is $\hat{u}_{k}^{c}$ if $\phi_{k}=1$, or the previous input $u_{k-1}$ if $\phi_{k}=0$, i.e., a zero-order hold mechanism is used in the event of downlink data loss (cf. (3)). The controller evolves based on an internal state $\hat{x}_{k}^{c}$, representing the estimate of plant state at time $k$ based on the available information and applies a state-feedback control input $\mathcal{F}\hat{x}_{k}^{c}$ (cf. (4) and (8)).
@@
Note that the controller does not have access to the variables $\phi_{k}$ in (6) since these occur on the actuator side and no feedback on $\phi_{k}$ is provided to the controller. Consequently, when constructing the networked control system dt-SLS-$\Sigma$ in Definition 1, the controller always assumes that the control input applied to the plant $\Sigma_{P}$ is $\hat{u}_{k}^{c}$ (i.e., $\phi_{k}=1$). Then, we define the controller’s state at time $k$ as $\hat{x}^{c}_{k}=\mathds{E}\big{[}{x}_{k}\,\big{|}\,\phi_{k}=1,\forall\,k\,,\mathcal{I}_{k}\big{]}$, where $\mathcal{I}_{k}$ is the set of available information at the control side at time $k$, defined as
$\mathcal{I}_{k}\;=\;\Bigl{\{}$ $A,B,x_{0},\,\{\theta_{0},\dots,\theta_{k-\tau}\},$
@@
$+\sum_{t=0}^{\tau-1}A^{t}w_{k-t-1}\Big{)}\!+\underbrace{(1-\theta_{k-\tau})(A\hat{x}^{c}_{k-1}\!+\!B\hat{u}_{k-1}^{c})}_{(\star)}\,\big{|}\,\mathcal{I}_{k}\Big{]}\!.$
We have formulated $(\star)$ based on the information set $\mathcal{I}_{k}$ available to the control unit at time $k$. If $\theta_{k-\tau}=0$, the control unit does not receive the plant state $x_{k-\tau}$, which is the most recent state the controller could access at time $k$ due to the $\tau$-step delay. In this case, the control unit uses the model in (4) to compute a model-based estimate of the state $\hat{x}_{k}^{c}$, which serves as the state of the controller. Since $w_{k}$ is a sequence of i.i.d. random variables drawn from $\mathcal{N}(0,\sigma_{w})$, and $\mathds{E}[\theta_{k-\tau}]=p_{\theta}$, we can rewrite (7) as
@@
Given a dt-SLS-$\Sigma=(X,U,A,B,w,\tau,\theta,\phi)$  as in Definition 1 subjected to delay and data loss, consider a safety specification $\Upsilon=(X_{0},X_{1},\mathcal{T})$, where $X_{0},X_{1}\subseteq X$ are its initial and unsafe sets, respectively. We assume $X_{0}\cap X_{1}=\emptyset$, otherwise system is unsafe with a probability of 1. The dt-SLS-$\Sigma$ is said to be safe within time horizon $\mathcal{T}\in\mathbb{N}$, denoted by $\Sigma\models\Upsilon$, if all trajectories of $\Sigma$ starting from $X_{0}$ do not reach $X_{1}$ during $\mathcal{T}$. Since trajectories of dt-SLS-$\Sigma$ are probabilistic, we are interested in computing $\mathds{P}\{\Sigma\models\Upsilon\}\geq 1-\xi_{\tau\theta\phi}$, (cf. Remark 2) where $\xi_{\tau\theta\phi}\in(0,1]$.